### v1

# strong corks derived from the Akbulut cork

Abstract.
Using instanton theory, we prove that the boundaries of the corks introduced by Auckly, Kim, Melvin, and Ruberman in [AKMR14] are strong corks, and that any nontrivial linear combination of them also yields a strong cork. Furthermore, via Heegaard Floer theory, we show that they are strong corks with respect to a distinct involution, and we construct a larger family of strong corks that generalizes them. We also prove that the boundaries of the corks introduced by Tange [Tan16] are strong corks and that any nontrivial linear combination of them yields a strong cork.
1. Introduction
In [Akb91], Akbulut provided the first example of a corkcork, which is defined as a compact, contractible 4-manifold $W$ equipped with an orientation-preserving diffeomorphism $\tau$ on its boundary that does not extend to a diffeomorphism of WW. [^1] Some authors require $W$ to be Stein, such as in [AY08]. By Freedman’s theorem [Fre82], $\tau$ extends over $W$ as a homeomorphism. Akbulut proved that the smooth structure of $K3\#\overline{\mathbb{C}P^{2}}$ can be changed by cutting out an embedded copy of $W$ and regluing it via $\tau$. Such an operation is called a corktwistcork\ twist. More generally, Matveyev [Mat95] and Curtis-Freedman-Hsiang-Stong [CFcHS96] showed that any two exotic smooth structures on a closed, simply-connected 4-manifold are related by a cork twist. Numerous examples of corks have been constructed; see, for example, [AY08] and [Gom17]. Corks are typically detected by embedding them into a closed 4-manifold with a non-vanishing smooth 4-manifold invariant, or by utilizing Stein structures and the adjunction inequality [AM97].
Recently, Lin, Ruberman, and Saveliev introduced the notion of a strongcorkstrong\ cork in [LRS18]. A strong cork is defined as a homology 3-sphere $Y$ which bounds at least one compact, contractible 4-manifold, equipped with an orientation-preserving involution $\tau$ that does not extend as a diffeomorphism over anyany smooth homology 4-ball bounded by $Y$. In [LRS18], they proved that the boundary of the Akbulut cork is a strong cork using monopole Floer homology. Subsequently, in [DHM20], Dai, Hedden, and Mallick exhibited many other examples of strong corks via Heegaard Floer theory. Some of them were later used to construct new closed exotic 4-manifolds in [LLP23]. Further methods for detecting strong corks have also been developed using family Seiberg-Witten theory [KMT23] and instanton theory [ADMT23].
Given a cork, we get an absolutely exotic pair of 4-manifolds as a consequence of the work of Akbulut-Ruberman [AR16]. In particular, given a strong cork $(Y,\tau)$, there exists a homology cobordism $W$ from $Y$ to another homology 3-sphere $Y^{\prime}$ such that, for anyany compact contractible 4-manifold $X$ bounded by $Y$, $X\cup_{\textrm{id}}W$ and $X\cup_{\tau}W$ form an exotic pair. Note that $X\cup_{\textrm{id}}W$ is homotopy equivalent to $X$, see also [Yas25]. This is one of the reasons why strong corks are important in the study of exotic 4-manifolds. In [Kan22], Kang constructed a cork $(C,\tau)$ such that $\tau$ does not extend over $C\#(S^{2}\times S^{2})$ as a diffeomorphism. Combined with the consequence in [AR16], it is revealed that there exists an absolutely exotic pair that remains absolutely exotic after one stabilization. See also [GK24]. In addition, (strong) corks have applications in the study of exotic disks and exotic surfaces; see, for example, [Akb22, Hay20, DMS22, HKM23, KMT24].
In light of the connections with exotic phenomena mentioned above, it is important to determine whether a cork is strong. We pose the following question:
Question 1.1.
Let $(C,\tau)$ be a cork. Is the boundary $(\partial C,\tau)$ a strong cork?
It has been shown that many corks, including Akbulut-Yasui corks $(W_{n},\tau)$ in [AY08], the first member $(\overline{W}_{1},\tau)$ of the positron corks, are strong. On the other hand, examples of non-strong corks were also constructed in [HP20]. The following are examples of known order 2 corks for which it has not been explicitly stated whether they are strong:
  1. (1)
    The positron cork $(\overline{W}_{n},\tau)$ for $n\geq 2$. This family was introduced by Akbulut-Matveyev in [AM00, Figure 2], where $(\overline{W}_{1},\tau)$ is the first positron mentioned above. See also [AY08, Figure 1].
  2. (2)
    The corks $(C_{n},\tau)$ and $(\overline{C}_{n},\tau)$, displayed in Figure 1, for even integers $n\geq 2$. These families were introduced by Auckly-Kim-Melvin-Ruberman in [AKMR14, Figure 8]. Note that $(C_{1},\tau)$ corresponds to the Akbulut cork $(W_{1},\tau)$, and $(\overline{C}_{1},\tau)$ corresponds to the first positron $(\overline{W}_{1},\tau)$. For odd $n$, it is shown in [DHM20] that $(\partial C_{n},\tau)$ and $(\partial\overline{C}_{n},\tau)$ are strong corks. The family of corks $(C_{n},\tau)$ was used to produce infinitely many exotic pairs of 4-manifolds in [Yas15].
  3. (3)
    The corks $(C(m),\tau)$, displayed in Figure 2, for $m\geq 2$. This family was introduced by Tange in [Tan16, Figure 1]. Note that $(C(1),\tau)$ corresponds to the Akbulut cork. This family of corks $(C(m),\tau)$ was used to produce many examples of finite order corks in [Tan16].
Figure 1. Left: $(C_{n},\tau)$. Here, the box labeled $\dfrac{-n+1}{2}$ represents n1n-1 negative half twists. Right: $(\overline{C}_{n},\tau)$.
Figure 2. $(C(m),\tau)$.
To the best of the author’s knowledge, it is unknown whether the boundaries of these corks are strong. In this paper, we address this question for specific families of corks. The families of strong corks constructed in this paper are derived from the Akbulut cork; the first member of each family is given by the boundary of the Akbulut cork.
1.1. Main results
First, we construct an infinite family of strong corks that generalizes the aforementioned examples of corks $(C_{n},\tau)$ and $(C(m),\tau)$:
Theorem 1.2.
For any $m,n\in\mathbb{N}$, let $(Z_{m,n},\tau)$ be the family of two-component link surgeries displayed in Figure 3. Equip $Z_{m,n}$ with the indicated involution $\tau$. Then $(Z_{m,n},\tau)$ is a strong cork. Furthermore, any nontrivial linear combination of elements in either $\{(Z_{m,n},\tau)\}_{m\in\mathbb{N}}$ (fixing $n$) or $\{(Z_{m,n},\tau)\}_{n\in\mathbb{N}}$ (fixing $m$) yields a strong cork. Specifically, for any sequence of integers $(a_{1},a_{2},\cdots,a_{k})\neq(0,0,\cdots,0)$, the equivariant connected sums
$(a_{1}Z_{1,n}\#a_{2}Z_{2,n}\#\cdots\#a_{k}Z_{k,n},\tau)\ and\ \ (a_{1}Z_{m,1}\#a_{2}Z_{m,2}\#\cdots\#a_{k}Z_{m,k},\tau)$
are strong corks. Here, for $a<0$, the notation $aZ$ denotes |a|(Z)|a|(-Z).
Figure 3. $(Z_{m,n},\tau)$
For $m=1$ and odd $n$, it is shown in [DHM20, Theorem 1.12] and [ADMT23, Theorem 1.5] that $(Z_{1,n},\tau)$ is a strong cork. However, linear combinations of these corks are not discussed in these papers.
When $m=1$, $(Z_{1,n},\tau)$ coincides with the boundary $(\partial C_{n},\tau)$ of Auckly-Kim-Melvin-Ruberman’s cork, and when $n=1$, $(Z_{m,1},\tau)$ coincides with the boundary $(\partial C(m),\tau)$ of Tange’s cork. In particular, for $m=n=1$, $(Z_{1,1},\tau)$ corresponds to the boundary of the Akbulut cork.
The proof of Theorem 1.2 relies on the instanton-theoretic invariant $r_{s}(Y,\tau)$ introduced in [ADMT23].
On the other hand, using the Heegaard-Floer-theoretic invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$ introduced in [DHM20], we construct the family of strong corks in the following theorem:
Theorem 1.3.
For $m,n\in\mathbb{N}$, let $Y_{m,n}$ be the family of two-component link surgeries displayed in Figure 4. Equip $Y_{m,n}$ with the indicated involution $\sigma$. Then $(Y_{m,n},\sigma)$ is a strong cork. It is also a strong cork by introducing any number of symmetric pairs of negative full twists.
Figure 4. $(Y_{m,n},\sigma)$.
The proof of Theorem 1.3 relies on the non-triviality of $h_{\iota\circ\sigma}(\Sigma(2,2m+1,4m+3))$, where $\Sigma(2,2m+1,4m+3)$ is a Brieskorn homology sphere equipped with the involution $\sigma$ displayed in Figure 7.
When $n=1$, $Y_{m,1}$ is diffeomorphic to the boundary $\partial W_{m}$ of the Akbulut-Yasui cork in [AY08, Figure 1]. For the case $m=1$, $(Y_{1,n},\sigma)$ is shown in Figure 5. The following theorem can be proved analogously to Theorem 1.2:
Theorem 1.4.
Any nontrivial linear combination of elements in $\{(Y_{1,n},\sigma)\}_{n\in\mathbb{N}}$ is a strong cork.
Figure 5. $(Y_{1,n},\sigma)$.
As shown in [AKMR14, Theorem 4.2], $(Z_{1,n},\tau)$, appearing in Theorem 1.2, is equivariantly diffeomorphic to $(S^{3}_{+1}(K_{n}),\tau)$, where $(K_{n},\tau)$ is the strongly invertible slice knot displayed on the left side of Figure 6. Using a technique similar to that in [AKMR14, Theorem 4.2], one can show that $(Y_{1,n},\sigma)$ is equivariantly diffeomorphic to $(S^{3}_{+1}(K_{n}),\sigma)$, where $(K_{n},\sigma)$ is displayed on the right side of Figure 6. The following theorem is an extension of [ADMT23, Theorem 1.5]:
Theorem 1.5.
For $n\in\mathbb{N}$, let $K_{n}$ be the strongly invertible slice knot displayed in Figure 6, equipped with the indicated involutions $\tau$ and $\sigma$. Then any nontrivial linear combination of elements in $\{(S^{3}_{1/m}(K_{n}),\tau)\}_{m\in\mathbb{N}}$ is a strong cork. This statement holds with $\tau$ replaced by $\sigma$.
Figure 6. Left: $(K_{n},\tau)$. Right: $(K_{n},\sigma)$.
When $n=1$, the first member of the family $K_{n}$ is $K_{1}=\overline{9}_{46}$, and $(S^{3}_{+1}(K_{1}),\tau)$ is the boundary of the Akbulut cork. In the case where $n$ is odd and the involution is $\tau$, Theorem 1.5 was established in [ADMT23].
The following proposition implies that the involutions of $(Z_{1,n},\tau)$ and $(Y_{1,n},\sigma)$ are induced by distinct symmetries of $K_{n}$, respectively, although $Z_{1,n}$ and $Y_{1,n}$ are diffeomorphic.
Proposition 1.6.
For any $n\in\mathbb{N}$, the knot $K_{n}$, displayed in Figure 6, admits two strong involutions $\tau$ and $\sigma$ such that the pairs $(K_{n},\tau)$ and $(K_{n},\sigma)$ are not Sakuma equivalent.
Using SnapPy [CDGW], we verified that, for $1\leq n\leq 5$, $K_{n}$ is a hyperbolic knot and the isometry group of $S^{3}\setminus K_{n}$ is isomorphic to $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$. By Mostow rigidity theorem [Mos68], if $K$ is a hyperbolic knot, then the symmetry group of $K$ is isomorphic to the isometry group of $S^{3}\setminus K$. The knot $K_{n}$ admits distinct strong involutions $\tau$ and $\sigma$. Since the product of two distinct strong involutions is orientation-preserving, the third non-trivial element of the group must be a periodic involution. Consequently, $K_{n}$ admits exactly two strong involutions for $1\leq n\leq 5$.
Question 1.7.
For any $n\geq 1$, is the knot $K_{n}$ hyperbolic? When $K_{n}$ is hyperbolic, is the isometry group of $S^{3}\setminus K_{n}$ isomorphic to $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$?
We prove Theorems 1.2, 1.3, 1.4, and 1.5 in Section 3, and we prove Proposition 1.6 in Section 4 (appendix).
We close this section with the following question.
Question 1.8.
Are the boundaries of the positron corks $(\overline{W}_{n},\tau)$ for $n\geq 2$ and Auckly-Kim-Melvin-Ruberman’s corks $(\overline{C}_{n},\tau)$ for even $n\geq 2$ (which are not covered in this paper) strong corks?
Acknowledgements
The author would like to express his gratitude to his advisor Masaki Taniguchi for helpful advice and encouragement. The author also would like to thank Irving Dai, Abhishek Mallick, Motoo Tange, and Kouichi Yasui for their valuable comments on the draft.
2. Definitions and background
2.1. Equivariant knots and equivariant surgery
We begin with the definition of an equivariant knot.
Definition 2.1.
Let $K$ be a knot in a 3-manifold $Y$, and let $\tau$ be an orientation-preserving involution on $Y$. We say that $(K,\tau)$ is an equivariantknotequivariant\ knot if $\tau$ fixes $K$ setwise. If $\tau$ fixes two points on $K$, we say that $(K,\tau)$ is stronglyinvertiblestrongly\ invertible. If $\tau$ has no fixed points on $K$, we say that $(K,\tau)$ is periodicperiodic.
Definition 2.2.
Two strongly invertible knots $(K,\tau)$ and $(K^{\prime},\tau^{\prime})$ are said to be SakumaequivalentSakuma\ equivalent if there exists an orientation-preserving diffeomorphism $f:S^{3}\rightarrow S^{3}$ such that $f(K)=K^{\prime}$ and $f\circ\tau=\tau^{\prime}\circ f$.
By the work of Waldhausen [Wal69], any orientation-preserving involution on $S^{3}$ is conjugate to a rotation about an unknot. In the case that $Y=S^{3}$, we will identify $S^{3}$ with the one-point compactification of $\mathbb{R}^{3}$ and often depict $\tau$ as a $180^{\circ}$ rotation around an axis in $\mathbb{R}^{3}$.
If two strongly invertible knots $(K,\tau)$ and $(K^{\prime},\tau^{\prime})$ are Sakuma equivalent, then they are related by a sequence of the involutive Reidemeister moves, together with an equivariant isotopy passing through the point at $\infty$ and a $180^{\circ}$ rotation about the origin; see [LW19, Theorem 2.4] and [BDMS25, Theorem 2.10].
In [DHM20, Section 5], it is shown that if $K$ is an equivariant knot in $Y$ with symmetry $\tau$, then $\tau$ extends to an involution on any Dehn surgery along $K$. We also denote it by $\tau$. Moreover, $\tau$ extends over the 2-handle cobordism given by attaching a 2-handle along $K$.
Similarly, if $L$ is a link equipped with $\tau$ that exchanges some pairs of components with the same framings, then $\tau$ extends to involutions on surgered manifolds and 2-handle cobordisms.
2.2. Strong corks
We recall the definitions of a cork and a strong cork.
Definition 2.3.
Let $C$ be a compact, contractible 4-manifold and $\tau:\partial C\rightarrow\partial C$ be an orientation-preserving diffeomorphism. We say that $(C,\tau)$ is a corkcork if $\tau$ does not extend over $C$ as a diffeomorphism.
In this paper, we do not require $C$ to be Stein, although this condition is sometimes required, such as in [AY08].
Definition 2.4.
Let $Y$ be an oriented integer homology 3-sphere and $\tau:Y\rightarrow Y$ be an orientation-preserving diffeomorphism on $Y$. Suppose that $Y$ bounds at least one compact, contractible manifold. We say that $(Y,\tau)$ is a $\mathit{strong\ cork}$ if $\tau$ does not extend as a diffeomorphism over any smooth homology ball that $Y$ bounds.
In this paper, we consider only the case where $\tau$ is an involution.
Definition 2.5.
[…]
If $(Y_{1},\tau_{1})$ and $(Y_{2},\tau_{2})$ are oriented homology 3-spheres equipped with orientation-preserving involutions, and the fixed-point set of $\tau_{i}$ is diffeomorphic to $S^{1}$, then we can define their equivariant connected sum $(Y_{1}\#Y_{2},\tau_{1}\#\tau_{2})$ as follows. Let $p_{1}\in Y_{1}$ and $p_{2}\in Y_{2}$ be fixed points of $\tau_{1}$ and $\tau_{2}$, respectively. Let $f_{i}:D^{3}\rightarrow Y_{i}$ be an embedding satisfying the following conditions:
  1. (1)
    $f_{i}(0)=p_{i}$ and $\tau_{i}(f_{i}(D^{3}))=f_{i}(D^{3})$.
  2. (2)
    $f^{-1}_{i}\circ\tau_{i}\circ f_{i}(x,y,z)=(x,-y,-z)$.
  3. (3)
    The orientation of $\textrm{Fix}(\tau_{i})$ coincides with the orientation of $f_{i}(D^{3})\cap\textrm{Fix}(\tau_{i})$ induced from the natural orientation of $\{(x,0,0)\in D^{3})\}$.
We set $D^{3}_{1/2}=\{(x,y,z)\in D^{3}\mid x^{2}+y^{2}+z^{2}\leq 1/2\}$ and define $\phi:\partial D^{3}_{1/2}\rightarrow\partial D^{3}_{1/2}$ by $(x,y,z)\mapsto(x,-y,z)$. Then the connected sum is defined as $Y_{1}\#Y_{2}=(Y_{1}\setminus f_{1}(\textrm{Int}(D^{3}_{1/2})))\cup_{\phi}(Y_{2}\setminus f_{2}(\textrm{Int}(D^{3}_{1/2})))$ and we can define the involution $\tau_{1}\#\tau_{2}$ on $Y_{1}\#Y_{2}$ such that $(\tau_{1}\#\tau_{2})|_{Y_{i}\setminus f_{i}(D^{3}_{1/2})}=\tau_{i}$. This operation is independent of the choice of $p_{1}$ and $p_{2}$.
3. detecting strong corks
3.1. Heegaard-Floer-theoretic methods
A lot of new families of strong corks have been discovered using Heegaard-Floer-theoretic methods developed in [DHM20]. The main tool is the following theorem:
Theorem 3.1.
[DHM20, Theorem 1.1] Let $Y$ be an oriented integer homology 3-sphere and $\tau$ be an orientation-preserving involution on $Y$. Then there are two invariants
$h_{\tau}(Y)=[(CF^{-}(Y)[-2],\tau)],\ h_{\iota\circ\tau}(Y)=[(CF^{-}(Y)[-2],\iota\circ\tau)]$
associated with the pair $(Y,\tau)$. If either $h_{\tau}(Y)\neq 0$ or $h_{\iota\circ\tau}(Y)\neq 0$, then $\tau$ does not extend to a diffeomorphism of any homology ball bounded by $Y$. Here, we denote $0=h_{\textrm{id}}(S^{3})$.
The invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$ take values in $\mathfrak{I}$, the group of $\iota$-complexes modulo local equivalence, which was defined in [HMZ17]. We briefly review the definition of the invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$. In [DHM20, Section 4], it is shown that $\tau$ induces a homotopy involution $\tau:CF^{-}(Y)\rightarrow CF^{-}(Y)$ and $(CF^{-}(Y),\tau)$ constitutes an $\iota$-complex. Taking the local equivalence class of this $\iota$-complex gives
$h_{\tau}(Y)=[(CF^{-}(Y)[-2],\tau)]\in\mathfrak{I}.$
Considering the map $\iota\circ\tau$ instead of $\tau$ gives another $\iota$-complex and its local equivalence class
$h_{\iota\circ\tau}(Y)=[(CF^{-}(Y)[-2],\iota\circ\tau)]\in\mathfrak{I}.$
The local equivalence group $\mathfrak{I}$ admits a partial order; see [DHM20, Definition 3.6]. The invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$ are monotonic under an appropriate equivariant negative-definite cobordism, as in the following theorem:
Theorem 3.2.
[DHM20, Theorem 1.5] Let $(Y,\tau)$ and $(Y^{\prime},\tau^{\prime})$ be oriented integer homology 3-spheres equipped with orientation-preserving involutions $\tau$ and $\tau^{\prime}$.
  1. (1)
    Let $(K,\tau)$ be an equivariant knot in $Y$. Suppose that $Y^{\prime}$ is obtained from $Y$ by performing (1)(-1)-surgery on $K$, and $\tau^{\prime}$ is the extension of $\tau$. If $(K,\tau)$ is periodic, then
    $h_{\tau}(Y)\leq h_{\tau^{\prime}}(Y^{\prime}).$
    If $(K,\tau)$ is strongly invertible, then
    $h_{\iota\circ\tau}(Y)\leq h_{\iota\circ\tau^{\prime}}(Y^{\prime}).$
  2. (2)
    Let $L=L_{1}\cup L_{2}$ be a two-component link in $Y$ whose components $L_{1},\ L_{2}$ are interchanged by $\tau$. The linking number of $L_{1},\ L_{2}$ is zero, and of $L_{i}$ with each component of the diagram of $Y$ is zero. Suppose that $Y^{\prime}$ is obtained from $Y$ by performing (1)(-1)-surgery on each component of $L$, and $\tau^{\prime}$ is the extension of $\tau$. Then
    $h_{\tau}(Y)\leq h_{\tau^{\prime}}(Y^{\prime})\ and\ h_{\iota\circ\tau}(Y)\leq h_{\iota\circ\tau^{\prime}}(Y^{\prime}).$
Attaching a 2-handle along $K$ (or 2-handles along $L$), as described in Theorem 3.2, yields an equivariant negative-definite cobordism from $(Y,\tau)$ to $(Y^{\prime},\tau^{\prime})$. In case (1) of Theorem 3.2, the resulting cobordism is called a $spin^{c}$-fixingcobordismfixing\ cobordism if $K$ is periodic, and is called a $spin^{c}$-conjugatingcobordismconjugating\ cobordism if $K$ is strongly invertible. In case (2), the corresponding cobordism is called an interchanging(1,1)interchanging\ (-1,-1)-cobordismcobordism.
Theorem 3.2 enables us to show the non-triviality of $h_{\tau}(Y)$ or $h_{\iota\circ\tau}(Y)$ by constructing an equivariant negative-definite cobordism from $(Y,\tau)$ to another pair $(Y^{\prime},\tau^{\prime})$ whose invariants are already understood. The following result is useful for constructing examples of strong corks:
Lemma 3.3.
[DHM20, Lemma 7.8] Let $T_{2,2n+1}$ be the right-handed $(2,2n+1)$-torus knot, equipped with the involutions $\tau$ and $\sigma$ indicated in Figure 7. Let $B_{n}=S^{3}_{-1}(T_{2,2n+1})=\Sigma(2,2n+1,4n+3)$. Then
$h_{\tau}(B_{n})=h_{\iota\circ\sigma}(B_{n})<0.$
Figure 7. $B_{n}=S^{3}_{-1}(T_{2,2n+1})=\Sigma(2,2n+1,4n+3)$
As shown in Figure 8, the knot $\overline{9}_{46}=P(-3,3,-3)$ admits two strong involutions $\tau$ and $\sigma$. It is known that $(S^{3}_{+1}(\overline{9}_{46}),\tau)$ is the boundary of the Akbulut cork. Note that $(Y_{1,1},\sigma)$ appearing in Theorem 1.3 is equivariantly diffeomorphic to $(S^{3}_{+1}(\overline{9}_{46}),\sigma)$.
Theorem 3.4.
[DHM20, Theorem 1.11] Let $Y=S^{3}_{+1}(\overline{9}_{46})$ be given by $(+1)$-surgery on the knot $\overline{9}_{46}$ equipped with the indicated involutions $\tau$ and $\sigma$. Then
$h_{\tau}(Y)<0,\ h_{\iota\circ\sigma}(Y)<0.$
Figure 8. $\overline{9}_{46}=P(-3,3,-3)$.
3.2. Proof of Theorem 1.3
The following lemma provides useful equivariant operations that will be used in the constructions below. These operations were used in [AKMR14, Theorem 4.2] without proof. Here, we show them explicitly.
Lemma 3.5.
The operations displayed in Figure 9 and 10 can be performed equivariantly. If the left and right components are distinct, their framings remain unchanged.
Figure 9. The operation in Lemma 3.5, where the central unknot has framing 1-1.
Figure 10. The operation in Lemma 3.5, where the central unknot has framing +1+1.
Proof.
We check the case where the central unknot has framing 1-1. See Figure 11 and 12. A similar deformation holds for the case where the central unknot has framing +1+1. ∎
Figure 11. The case where the central linking between the red and black components is positive.
Figure 12. The case where the central linking between the red and black components is negative.
Proof of Theorem 1.3.
We construct an interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)=(\Sigma(2,2m+1,4m+3),\sigma)$ as shown in Figure 13.222It can be verified that $Y_{m,1}$ is diffeomorphic to the boundary $\partial W_{m}$, where $(W_{m},\tau)$ is the Akbulut-Yasui cork in [AY08]. In [DHM20, Theorem 1.13], it is shown that $h_{\tau}(\partial W_{m})\leq h_{\tau}(B_{m})<0$, which implies that $(\partial W_{m},\tau)$ is a strong cork. This is formed by attaching (1)(-1)-framed 2-handles to the green unknots as in the top left of Figure 13. We see that the condition of Theorem 3.2 (2) are satisfied by simultaneously sliding the left green unknot over the right component of $Y_{m,1}$ and the right green unknot over the left component of $Y_{m,1}$. By Theorem 3.2 (2) and Lemma 3.3,
$h_{\iota\circ\sigma}(Y_{m,1})\leq h_{\iota\circ\sigma}(B_{m})<0.$
By Lemma 3.5, we obtain a spinc-conjugating (1)(-1)-cobordism from $(Y_{m,n+1},\sigma)$ to $(Y_{m,n},\sigma)$ as shown in Figure 14. By Theorem 3.2 (1),
$h_{\iota\circ\sigma}(Y_{m,n})\leq h_{\iota\circ\sigma}(Y_{m,1})<0.$
Therefore, it follows from Theorem 3.1 that $Y_{m,n}$ is a strong cork.
Figure 13. The interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)$ used in the proof of Theorem 1.3
Moreover, if $Y^{\prime}_{m,n}$ is constructed from $Y_{m,n}$ by introducing any number of symmetric pairs of negative full twists, then $Y^{\prime}_{m,n}$ admits a sequence of interchanging (1,1)(-1,-1)-cobordisms to $Y_{m,n}$. Thus $Y^{\prime}_{m,n}$ is also a strong cork. ∎
Figure 14. The spinc-conjugating (1)(-1)-cobordism from $(Y_{m,n+1},\sigma)$ to $(Y_{m,n},\sigma)$ used in the proof of Theorem 1.3
3.3. Instanton-theoretic methods
Instanton-theoretic methods developed in [ADMT23] enable us to detect strong corks and treat examples inaccessible to Heegaard Floer theory.
Theorem 3.6.
[ADMT23, Theorem 1.1] Let $(Y,\tau)$ be an oriented integer homology 3-sphere equipped with an orientation-preserving involution on $Y$. For $s\in[-\infty,0]$, there is a real number
$r_{s}(Y,\tau)\in(0,\infty]$
[…]
$r_{s}(Y,\tau)\leq r_{s}(Y^{\prime},\tau).$
If $r_{s}(Y,\tau)<\infty$ and $W$ is simply connected, then
$r_{s}(Y,\tau)<r_{s}(Y^{\prime},\tau).$
It follows from Theorem 3.6 that if $\tau$ extends over some homology ball $W$ that $Y$ bounds, then $r_{s}(Y,\tau)=\infty$ for any $s\in[-\infty,0]$. Thus, this invariant $r_{s}$ can be used to detect strong corks.
In contrast to Theorem 3.2 in Heegaard Floer theory, the inequality for the $r_{s}$-invariant holds regardless of whether the cobordism is spinc-fixing or spinc-conjugating, as in Theorem 3.6.
The following theorem allows us to treat linear combinations using the $r_{s}$-invariant.
Theorem 3.7.
[ADMT23, Theorem 7.7] Let $\{(Y_{i},\tau_{i})\}^{\infty}_{i=1}$ be a sequence of oriented integer homology 3-spheres equipped with orientation-preserving involutions $\tau_{i}$. Assume the fixed-point set of each $\tau_{i}$ is a copy of $S^{1}$, so that the equivariant connected sum operation is well-defined. Suppose that:
  1. (1)
    $r_{0}(Y_{1},\tau_{1})>r_{0}(Y_{2},\tau_{2})>\cdots>r_{0}(Y_{i},\tau_{i})>\cdots$,
  2. (2)
    $r_{0}(Y_{1},\tau_{1})<\infty$,
  3. (3)
    $r_{0}(-Y_{i},\tau_{i})=\infty$ for each $i$.
Then, any nontrivial linear combination of elements in $\{(Y_{i},\tau_{i})\}^{\infty}_{i=1}$ is a strong cork.
As stated in the following lemma, the boundary of the Akbulut cork is shown to be non-trivial with respect to the $r_{0}$-invariant:
Lemma 3.8.
[ADMT23, Lemma 7.1] Let $Y=S^{3}_{+1}(\overline{9}_{46})$ equipped with the indicated involutions $\tau$ and $\sigma$ displayed in Figure 8. Then
$r_{0}(Y,\tau)<\infty,\ r_{s}(-Y,\tau)=\infty.$
for any $s\in[-\infty,0]$. This statement holds with $\tau$ replaced by $\sigma$.
3.4. Proof of Theorems 1.2, 1.5, and 1.4
Proof of Theorem 1.2.
Using the operation in Lemma 3.5, we obtain an equivariant negative-definite cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$, as shown in Figure 15. Similarly, we obtain an equivariant negative-definite cobordism from $(Z_{m+1,n},\tau)$ to $(Z_{m,n},\tau)$. Since $(Z_{1,1},\tau)$ is equivariantly diffeomorphic to $(S^{3}_{+1}(\overline{9}_{46}),\tau)$, by Theorem 3.6 and Lemma 3.8,
$r_{0}(Z_{m,n},\tau)\leq r_{0}(Z_{1,1},\tau)<\infty.$
Therefore, for any $m,n\geq 1$, $(Z_{m,n},\tau)$ is a strong cork.
Figure 15. The equivariant negative-definite cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$ used in the proof of Theorem 1.2.
Claim 3.9.
This cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$ is simply connected.
Figure 16. The cobordism $W_{1,1}$ and the Wirtinger presentation of the surgery link for $Z_{1,2}.$
Proof.
First, we verify the case where $m,n=1$. The generators of $\pi_{1}(W_{1,1})$ are $x_{1},\cdots,x_{5},y_{1},\cdots,y_{5}$ as indicated in Figure 16. The relators are given by:
(1) $y^{-1}_{1}x^{-1}_{1}y_{5}x_{1}=1$
(2) $x^{-1}_{1}y^{-1}_{5}x_{5}y_{5}=1$
(3) $x^{-1}_{2}x^{-1}_{5}x_{1}x_{5}=1$
(4) $y^{-1}_{5}y^{-1}_{1}y_{4}y_{1}=1$
(5) $y^{-1}_{2}x^{-1}_{1}y_{1}x_{2}=1$
(6) $x^{-1}_{3}y^{-1}_{2}x_{2}y_{2}=1$
(7) $x^{-1}_{5}x^{-1}_{3}x_{4}x_{3}=1$
(8) $y^{-1}_{3}y^{-1}_{4}y_{2}y_{4}=1$
(9) $y^{-1}_{4}x^{-1}_{4}y_{4}x_{3}=1$
(10) $x^{-1}_{3}y^{-1}_{4}x_{3}y_{3}=1$
(11) $y_{5}x_{5}y_{2}x^{-2}_{3}y^{-1}_{4}x_{3}=1$
(12) $x_{1}x_{2}y_{4}y^{-2}_{3}x^{-1}_{3}y_{1}=1$
(13) $x_{1}x^{-1}_{5}=1$
The equations (1)-(10) correspond to the relators for the fundamental group of the complement of the surgery link for $Z_{1,2}$. The equations (11) and (12) correspond to the 0-framed longitudes of the components of the diagram of $Z_{1,2}$, respectively. The equation (13) corresponds to the green attaching sphere of the 2-handle of $W_{1,1}$.
From (13), we have $x_{1}=x_{5}$, and let us denote this element by $x$. If we set $y=y_{5}$, then (2) implies $xy=yx$. Thus, from (1), we have $y_{1}=y$. By (3), $x_{2}=x$, and by (4), $y_{4}=y$. By (5), $y_{2}=y$, so by (6), $x_{3}=x$. By (7), $x_{4}=x$, and by (8), $y_{3}=y$. In summary, from (1)-(8) and (13), we obtain
(14) $x=x_{1}=x_{2}=x_{3}=x_{4}=x_{5},\ y=y_{1}=y_{2}=y_{3}=y_{4}=y_{5},\ xy=yx.$
Substituting these into (11) and (12), we obtain
$x=y=1.$
Therefore, $W_{1,1}$ is simply connected.
Note that since $Z_{1,2}$ is an integer homology 3-sphere, its first homology group is trivial. Since the first homology group is the abelianization of the fundamental group, the relations in (14) imply $x=y=1$ immediately, without the need to substitute into (11) and (12).
Even when additional half twists are added, we can compute inductively in a similar way. Consider the case where a half twist is added as shown in Figure 17. For the left case, if we assume $x=x_{k+1}=x_{l}$, then the relator $x^{-1}_{k+1}x^{-1}_{l}x_{k}x_{l}=1$ implies $x=x_{k}$. Similarly, for the right case, assuming $x=x_{k}=x_{l}$, the relator $x^{-1}_{k+1}x^{-1}_{l}x_{k}x_{l}=1$ implies $x=x_{k+1}$. Thus, it is shown inductively that $W_{m,n}$ is simply connected for any $m$ and $n$. ∎
Figure 17.
By Claim 3.9, we have $r_{0}(Z_{m,{n+1}},\tau)<r_{0}(Z_{m,n},\tau)$. Considering $W_{m,n}$ to be the cobordism from $(-Z_{m,{n}},\tau)$ to $(-Z_{m,{n+1}},\tau)$, we obtain
$\infty=r_{0}(-Z_{1,1},\tau)\leq r_{0}(-Z_{m,1},\tau)\leq r_{0}(-Z_{m,2},\tau)\leq\cdots\leq r_{0}(-Z_{m,n},\tau)\leq\cdots.$
Thus, $r_{0}(-Z_{m,n},\tau)=\infty$. Therefore, any nontrivial linear combination of elements in $\{(Z_{m,n},\tau)\}_{n\in\mathbb{N}}$ is a strong cork by Theorem 3.7.
Similarly, the cobordism from $(Z_{m+1,n},\tau)$ to $(Z_{m,n},\tau)$ is also simply connected, so any nontrivial linear combination of elements in $\{(Z_{m,n},\tau)\}_{m\in\mathbb{N}}$ is a strong cork. ∎
Remark 3.10.
When $m$ and $n$ are odd, one can also prove that
$h_{\tau}(Z_{m,n})\leq h_{\tau}(Z_{1,1})<0$
and $(Z_{m,n},\tau)$ is strong by considering the interchanging (1,1)(-1,-1)-cobordism shown in Figure 18. However, since the cobordism used in the proof of Theorem 1.2 is spinc-conjugating, we cannot obtain the inequality for $h_{\tau}$. Thus, Theorem 1.2 cannot be immediately shown using Heegaard Floer theory when $m$ and $n$ are even.
Figure 18. The interchanging (1,1)(-1,-1)-cobordism from $(Z_{m,n+2},\tau)$ to $(Z_{m,n},\tau)$.
Proof of Theorem 1.5.
It suffices to prove that $\{(S^{3}_{1/m}(K_{n}),\tau)\}_{m\in\mathbb{N}}$ satisfies the assumptions of Theorem3.7.
  1. (1)
    The cobordism from $(S^{3}_{1/{m+1}}(K_{n}),\tau)$ to $(S^{3}_{1/m}(K_{n}),\tau)$ displayed in Figure 19 is simply-connected, equivariant negative-definite, which can be proved similarly to [NST19, Theorem 5.12].
  2. (2)
    Since $(S^{3}_{+1}(K_{n}),\tau)$ is equivariantly diffeomorphic to $(Z_{1,n},\tau)$, by Theorem 1.2,
    $r_{0}(S^{3}_{+1}(K_{n}),\tau)=r_{0}(Z_{1,n},\tau)<\infty.$
  3. (3)
    By [ADMT23, Theorem 6.3], $r_{0}(-S^{3}_{1/m}(K_{n}),\tau)=\infty$ for any $m\geq 1$.
The same holds for the involution $\sigma$. ∎
Figure 19. The simply-connected, equivariant negative-definite cobordism from $(S^{3}_{1/{m+1}}(K_{n}),\tau)$ to $(S^{3}_{1/m}(K_{n}),\tau)$ used in the proof of Theorem 1.5.
Remark 3.11.
Theorem 1.3 cannot immediately be proved by instanton theory using the cobordism in the proof of Theorem 1.3 because the $r_{s}$-invariants are trivial for Brieskorn spheres, as discussed in [ADMT23, Section 7.2].
Proof of Theorem 1.4.
Since $(Y_{1,n},\sigma)$ is equivariantly diffeomorphic to $(S^{3}_{+1}(\overline{9}_{46}),\sigma)$, by Lemma 3.8,
$r_{0}(Y_{1,n},\sigma)=r_{0}(S^{3}_{+1}(\overline{9}_{46}),\sigma)<\infty.$
Similarly to the proof of Theorem 1.2, there is a simply-connected, equivariant negative-definite cobordism from $(Y_{1,n+1},\sigma)$ to $(Y_{1,n},\sigma)$. The result follows from Theorem 3.7. ∎
4. appendix
In this appendix, we prove Proposition 1.6, which states that the knot $K_{n}$ has two distinct strong involutions $\tau$ and $\sigma$ displayed in Figure 6.
Lemma 4.1.
For each $n\in\mathbb{N}$, the knot displayed on the left side of Figure 6 is isotopic to the one displayed on the right.
Proof.
See Figure 26. ∎
It remains to prove that the involutions $\tau$ and $\sigma$ are distinct. We prove that $(K_{n},\tau)$ and $(K_{n},\sigma)$ are not Sakuma equivalent by computing their Sakuma $\eta$-polynomials introduced in [Sak86].
We review the definition of the $\eta$-polynomial of a strongly invertible knot $(K,\tau)$. Let $l$ be a preferred longitude of $K$ such that $\tau(l)\cap l=\emptyset$. Set $O=p(\textrm{Fix}(\tau))$ and $L=p(l)$, where $p:S^{3}\rightarrow S^{3}/\tau\cong S^{3}$ is a projection. Then $L(K,\tau)=O\cup L$ is a two-component link in $S^{3}$ and $\textrm{lk}(O,L)=0$. The $\eta$-polynomial of $(K,\tau)$ is defined as the Laurent polynomial
$\eta_{(K,\tau)}(t)=\sum_{i=-\infty}^{\infty}\textrm{lk}(\tilde{L^{\prime}},t^{i}\tilde{L})t^{i}\in\mathbb{Z}[t^{\pm 1}],$
where $\tilde{L}$ is the lift of $L$ to the infinite cyclic cover $\widetilde{E(O)}$ of $E(O)=S^{3}\setminus O$, $\tilde{L^{\prime}}$ is the lift of a preferred longitude $L^{\prime}$ of $L$ near $\tilde{L}$, and $t$ is a generator of the covering transformation group of $\widetilde{E(O)}$. The $\eta$-polynomial $\eta_{(K,\tau)}(t)$ is an invariant of the Sakuma equivalence class of $(K,\tau)$.
Lemma 4.2.
For each $n\in\mathbb{N}$, $\eta_{(K_{n},\tau)}(t)\neq\eta_{(K_{n},\sigma)}(t)$.
Figure 20. Left: $(K_{n},\tau)$. Right: The intermediate step to obtain the pseudo-fundamental region.
Figure 21. Left: The pseudo-fundamental region of $(K_{n},\tau)$ where $n$ is odd. Indices are assigned to the arcs according to Sakuma’s algorithm. Starting with 0, the index is increased by 1 if the arc starts from the right, and decreased by 1 if it starts from the left. At each crossing, red and blue circles indicate positive and negative signs, respectively, and the number represents the index of the over-arcs minus that of the under-arcs. Right: A detail of one of the full twists in the box on the left.
Proof.
We use the algorithm described in [Sak86, Section 2]. First, we compute the Sakuma $\eta$-polynomial $\eta_{(K_{n},\tau)}(t)$ of $(K_{n},\tau)$.
The case where $n$ is odd: Via the process in Figure 20, we obtain the ”pseudo-fundamental region” of $(K_{n},\tau)$ in Figure 21, which gives the calculation of the $\eta$-polynomial.
$\tilde{\eta}_{(K_{n},\tau)}=-\dfrac{n-1}{2}(x_{-3}+x_{3})+(-n+2)(x_{-2}+x_{2})+\dfrac{n+3}{2}(x_{-1}+x_{1})+2x_{0}.$
Putting $x_{i}=t^{i-1}-2t^{i}+t^{i+1}$,
$\eta^{\prime}_{(K_{n},\tau)}=(n+1)+(-2n-1)(t^{-1}+t)+(2n-2)(t^{-2}+t^{2})+(t^{-3}+t^{3})-\dfrac{n-1}{2}(t^{-4}+t^{4}).$
If $\eta^{\prime}=[a_{0},a_{1},a_{2},a_{3},\cdots$, then,
$\eta=[-2\Sigma_{j\geq 1}a_{2j},\ -\Sigma_{j\geq 1}a_{2j+1},\ a_{2},a_{3},\cdots.$
Here, $a_{0},a_{1},a_{2},\cdots,a_{n}$ represents the polynomial $a_{0}+a_{1}(t^{-1}+t)+a_{2}(t^{-2}+t^{2})+\cdots+a_{n}(t^{-n}+t^{n})$. Hence, the $\eta$-polynomial of $(K_{n},\tau)$ is
$\eta_{(K_{n},\tau)}=[-3n+3,\ -1,\ 2n-2,\ 1,\ -\dfrac{n-1}{2}.$
The case where $n$ is even: Figure 22 shows the pseudo-fundamental region of $(K_{n},\tau)$ where $n$ is even.
$\tilde{\eta}_{(K_{n},\tau)}=-\dfrac{n}{2}(x_{-5}+x_{5})+(n-1)(x_{-3}+x_{3})+(-n+1)(x_{-2}+x_{2})+\left(-\dfrac{n}{2}+1\right)(x_{-1}+x_{1}).$
Putting $x_{i}=t^{i-1}-2t^{i}+t^{i+1}$,
$\eta^{\prime}_{(K_{n},\tau)}=(-n+2)-(t^{-1}+t^{1})+\left(\dfrac{5}{2}n-2\right)(t^{-2}+t^{2})+(-3n+3)(t^{-3}+t^{3})+\left(\dfrac{n}{2}-1\right)(t^{-4}+t^{4})+n(t^{-5}+t^{5})-\dfrac{n}{2}(t^{-6}+t^{6}).$
Hence,
$\eta_{(K_{n},\tau)}=[-5n+6,\ 2n-3,\ \dfrac{5}{2}n-2,\ -3n+3,\ \dfrac{n}{2}-1,\ n,\ -\dfrac{n}{2}.$
Figure 22. Left: The pseudo-fundamental region of $(K_{n},\tau)$ where $n$ is even. Right: A detail of one of the full twists in the box on the left.
Next, we compute the $\eta$-polynomial $\eta_{(K_{n},\sigma)}(t)$ of $(K_{n},\sigma)$.
The case where $n$ is odd: Via the process in Figure 23, we obtain the pseudo-fundamental region of $(K_{n},\sigma)$ in Figure 24.
$\tilde{\eta}_{(K_{n},\sigma)}=\dfrac{n-1}{2}(x_{-5}+x_{5})+(-n+2)(x_{-3}+x_{3})+(n-2)(x_{-2}+x_{2})+\dfrac{n-3}{2}(x_{-1}+x_{1}).$
Putting $x_{i}=t^{i-1}-2t^{i}+t^{i+1}$,
$\eta^{\prime}_{(K_{n},\sigma)}=(n-3)+(t^{-1}+t)+\dfrac{-5n+9}{2}(t^{-2}+t^{2})+(3n-6)(t^{-3}+t^{3})+\dfrac{-n+3}{2}(t^{-4}+t^{4})+(-n+1)(t^{-5}+t^{5})+\dfrac{n-1}{2}(t^{-6}+t^{6}).$
Hence,
$\eta_{(K_{n},\sigma)}=[5n-11,\ -2n+5,\ \dfrac{-5n+9}{2},\ 3n-6,\ \dfrac{-n+3}{2},\ -n+1,\ \dfrac{n-1}{2}.$
Figure 23. Left: $(K_{n},\sigma)$. Right: The intermediate step to obtain the pseudo-fundamental region.
Figure 24. Left: The pseudo-fundamental region of $(K_{n},\sigma)$ where $n$ is odd. Right: A detail of one of the full twists in the box on the left.
The case where $n$ is even: Figure 25 shows the pseudo-fundamental region of $(K_{n},\sigma)$ where $n$ is even.
$\tilde{\eta}_{(K_{n},\sigma)}=\dfrac{n-4}{2}(x_{-3}+x_{3})+(n-5)(x_{-2}+x_{2})-\dfrac{n}{2}(x_{-1}+x_{1})-2x_{0}.$
Putting $x_{i}=t^{i-1}-2t^{i}+t^{i+1}$,
$\eta^{\prime}_{(K_{n},\sigma)}=\dfrac{n-4}{2}(t^{-4}+t^{4})-(t^{-3}+t^{3})+(-2n+8)(t^{-2}+t^{2})+(2n-7)(t^{-1}+t)+(-n+4).$
Hence,
$\eta_{(K_{n},\sigma)}=[3n-12,\ 1,\ -2n+8,\ -1,\ \dfrac{n-4}{2}.$
Figure 25. Left: The pseudo-fundamental region of $(K_{n},\sigma)$ where $n$ is even. Right: A detail of one of the full twists in the box on the left.
Figure 26. The isopoty of $K_{n}$ used in the proof of Lemma 4.1

### v2

@@
 Abstract.
-Using instanton theory, we prove that the boundaries of the corks introduced by Auckly, Kim, Melvin, and Ruberman in [AKMR14] are strong corks, and that any nontrivial linear combination of them also yields a strong cork.
-Furthermore, via Heegaard Floer theory, we show that they are strong corks with respect to a distinct involution, and we construct a larger family of strong corks that generalizes them.
-We also prove that the boundaries of the corks introduced by Tange [Tan16] are strong corks and that any nontrivial linear combination of them yields a strong cork.
+We prove that the boundaries of the corks introduced by Auckly, Kim, Melvin, and Ruberman in [AKMR14] and by Tange in [Tan16] are strong corks.
+Furthermore, we prove that any nontrivial linear combination of them yields a strong cork, and we construct a larger family of strong corks that generalizes them.
+These results rely on the instanton-theoretic invariant introduced by Alfieri, Dai, Mallick, and Taniguchi in [ADMT23].
 1. Introduction
-In [Akb91], Akbulut provided the first example of a corkcork, which is defined as a compact, contractible 4-manifold $W$ equipped with an orientation-preserving diffeomorphism $\tau$ on its boundary that does not extend to a diffeomorphism of WW.
-[^1] Some authors require $W$ to be Stein, such as in [AY08].
-By Freedman’s theorem [Fre82], $\tau$ extends over $W$ as a homeomorphism.
+In [Akb91], Akbulut provided the first example of a corkcork, which is defined as a compact, contractible 4-manifold $W$ equipped with an orientation-preserving diffeomorphism $\tau$ on its boundary that does not extend to a diffeomorphism of $W$.
+(Some authors require $W$ to be Stein, such as in [AY08].) By Freedman’s theorem [Fre82], $\tau$ extends over $W$ as a homeomorphism.
 Akbulut proved that the smooth structure of $K3\#\overline{\mathbb{C}P^{2}}$ can be changed by cutting out an embedded copy of $W$ and regluing it via $\tau$.
@@
 These families were introduced by Auckly-Kim-Melvin-Ruberman in [AKMR14, Figure 8].
-Note that $(C_{1},\tau)$ corresponds to the Akbulut cork $(W_{1},\tau)$, and $(\overline{C}_{1},\tau)$ corresponds to the first positron $(\overline{W}_{1},\tau)$.
+(For [AKMR14], see also [Akb15].) Note that $(C_{1},\tau)$ corresponds to the Akbulut cork $(W_{1},\tau)$, and $(\overline{C}_{1},\tau)$ corresponds to the first positron $(\overline{W}_{1},\tau)$.
 For odd $n$, it is shown in [DHM20] that $(\partial C_{n},\tau)$ and $(\partial\overline{C}_{n},\tau)$ are strong corks.
@@
 The proof of Theorem 1.2 relies on the instanton-theoretic invariant $r_{s}(Y,\tau)$ introduced in [ADMT23].
+Remark 1.3.
+$(Z_{m,n}\#(-Z_{m,n}),\tau)$ is a boundary of a cork which is not strong; that is, $\tau$ extends over some homology ball that $Z_{m,n}\#(-Z_{m,n})$ bounds.
+However, $\tau$ does not extend over any contractible 4-manifold that $Z_{m,n}\#(-Z_{m,n})$ bounds.
+See Remark 3.12.
 On the other hand, using the Heegaard-Floer-theoretic invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$ introduced in [DHM20], we construct the family of strong corks in the following theorem:
@@
 Consequently, $K_{n}$ admits exactly two strong involutions for $1\leq n\leq 5$.
-Question 1.7.
+Question 1.8.
 For any $n\geq 1$, is the knot $K_{n}$ hyperbolic? When $K_{n}$ is hyperbolic, is the isometry group of $S^{3}\setminus K_{n}$ isomorphic to $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$?
-We prove Theorems 1.2, 1.3, 1.4, and 1.5 in Section 3, and we prove Proposition 1.6 in Section 4 (appendix).
+We prove Theorems 1.2, 1.4, 1.5, and 1.6 in Section 3, and we prove Proposition 1.7 in Section 4 (appendix).
 We close this section with the following question.
-Question 1.8.
+Question 1.9.
 Are the boundaries of the positron corks $(\overline{W}_{n},\tau)$ for $n\geq 2$ and Auckly-Kim-Melvin-Ruberman’s corks $(\overline{C}_{n},\tau)$ for even $n\geq 2$ (which are not covered in this paper) strong corks?
@@
 Considering the map $\iota\circ\tau$ instead of $\tau$ gives another $\iota$-complex and its local equivalence class
-$h_{\iota\circ\tau}(Y)=[(CF^{-}(Y)[-2],\iota\circ\tau)]\in\mathfrak{I}.$
+$h_{\iota\circ\tau}(Y)=[(CF^{-}(Y)[-2],\iota\circ\tau)]\in\mathfrak{I},$
+where $\iota$ is the homotopy involution on $CF^{-}(Y)$ defined in [HM17].
 The local equivalence group $\mathfrak{I}$ admits a partial order; see [DHM20, Definition 3.6].
@@
 The following result is useful for constructing examples of strong corks:
+Figure 7. $B_{n}=S^{3}_{-1}(T_{2,2n+1})=\Sigma(2,2n+1,4n+3)$
 Lemma 3.3.
@@
 $h_{\tau}(B_{n})=h_{\iota\circ\sigma}(B_{n})<0.$
-Figure 7. $B_{n}=S^{3}_{-1}(T_{2,2n+1})=\Sigma(2,2n+1,4n+3)$
 As shown in Figure 8, the knot $\overline{9}_{46}=P(-3,3,-3)$ admits two strong involutions $\tau$ and $\sigma$.
@@
 Figure 8. $\overline{9}_{46}=P(-3,3,-3)$.
+Figure 9. The operation in Lemma 3.5, where the central unknot has framing 1-1.
+Figure 10. The operation in Lemma 3.5, where the central unknot has framing +1+1.
 3.2. Proof of Theorem 1.4
@@
 Here, we show them explicitly.
+Figure 11. The case where the central linking between the red and black components is positive.
+Figure 12. The case where the central linking between the red and black components is negative.
 Lemma 3.5.
 The operations displayed in Figure 9 and 10 can be performed equivariantly. If the left and right components are distinct, their framings remain unchanged.
-Figure 9. The operation in Lemma 3.5, where the central unknot has framing 1-1.
-Figure 10. The operation in Lemma 3.5, where the central unknot has framing +1+1.
 Proof.
 We check the case where the central unknot has framing 1-1. See Figure 11 and 12. A similar deformation holds for the case where the central unknot has framing +1+1. ∎
-Figure 11. The case where the central linking between the red and black components is positive.
-Figure 12. The case where the central linking between the red and black components is negative.
 Proof of Theorem 1.4.
-We construct an interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)=(\Sigma(2,2m+1,4m+3),\sigma)$ as shown in Figure 13.222It can be verified that $Y_{m,1}$ is diffeomorphic to the boundary $\partial W_{m}$, where $(W_{m},\tau)$ is the Akbulut-Yasui cork in [AY08].
-In [DHM20, Theorem 1.13], it is shown that $h_{\tau}(\partial W_{m})\leq h_{\tau}(B_{m})<0$, which implies that $(\partial W_{m},\tau)$ is a strong cork.
-This is formed by attaching (1)(-1)-framed 2-handles to the green unknots as in the top left of Figure 13.
+We construct an interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)=(\Sigma(2,2m+1,4m+3),\sigma)$ as shown in Figure 13.
+This is formed by attaching (1)(-1)-framed 2-handles along the green unknots as in the top left of Figure 13.
 We see that the condition of Theorem 3.2 (2) are satisfied by simultaneously sliding the left green unknot over the right component of $Y_{m,1}$ and the right green unknot over the left component of $Y_{m,1}$.
@@
 Therefore, it follows from Theorem 3.1 that $Y_{m,n}$ is a strong cork.
-Figure 13. The interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)$ used in the proof of Theorem 1.3
+Figure 13.
+The interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)$ used in the proof of Theorem 1.4, obtained by attaching (1)(-1)-framed 2-handles along the green unknots.
 Moreover, if $Y^{\prime}_{m,n}$ is constructed from $Y_{m,n}$ by introducing any number of symmetric pairs of negative full twists, then $Y^{\prime}_{m,n}$ admits a sequence of interchanging (1,1)(-1,-1)-cobordisms to $Y_{m,n}$.
@@
 ∎
-Figure 14. The spinc-conjugating (1)(-1)-cobordism from $(Y_{m,n+1},\sigma)$ to $(Y_{m,n},\sigma)$ used in the proof of Theorem 1.3
+Remark 3.6.
+It can be verified that $Y_{m,1}$ is diffeomorphic to the boundary $\partial W_{m}$, where $(W_{m},\tau)$ is the Akbulut-Yasui cork in [AY08].
+In [DHM20, Theorem 1.13], it is shown that $h_{\tau}(\partial W_{m})\leq h_{\tau}(B_{m})<0$, which implies that $(\partial W_{m},\tau)$ is a strong cork.
+Figure 14.
+The spinc-conjugating (1)(-1)-cobordism from $(Y_{m,n+1},\sigma)$ to $(Y_{m,n},\sigma)$ used in the proof of Theorem 1.4, obtained by attaching (1)(-1)-framed 2-handles along the green unknot
 3.3. Instanton-theoretic methods
@@
 [ADMT23, Theorem 7.7] Let $\{(Y_{i},\tau_{i})\}^{\infty}_{i=1}$ be a sequence of oriented integer homology 3-spheres equipped with orientation-preserving involutions $\tau_{i}$.
-Assume the fixed-point set of each $\tau_{i}$ is a copy of $S^{1}$, so that the equivariant connected sum operation is well-defined.
+Assume that each $Y_{i}$ bounds a compact, contractible 4-manifold, and that the fixed-point set of each $\tau_{i}$ is a copy of $S^{1}$, so that the equivariant connected sum operation is well-defined.
 Suppose that:
@@
 for any $s\in[-\infty,0]$. This statement holds with $\tau$ replaced by $\sigma$.
-3.4. Proof of Theorems 1.2, 1.5, and 1.4
+3.4. Proof of Theorems 1.2, 1.6, and 1.5
 Proof of Theorem 1.2.
@@
 Therefore, for any $m,n\geq 1$, $(Z_{m,n},\tau)$ is a strong cork.
-Figure 15. The equivariant negative-definite cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$ used in the proof of Theorem 1.2.
+Figure 15.
+The equivariant negative-definite cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$ used in the proof of Theorem 1.2, obtained by attaching (1)(-1)-framed 2-handles along the green unknot.
 Claim 3.10.
@@
 Figure 18. The interchanging (1,1)(-1,-1)-cobordism from $(Z_{m,n+2},\tau)$ to $(Z_{m,n},\tau)$.
+Remark 3.12.
+$(Z_{m,n}\#(-Z_{m,n}),\tau)$ is not a strong cork, but $\tau$ does not extend over any contractible 4-manifold bounded by $Z_{m,n}\#(-Z_{m,n})$.
+If $\tau$ were to extend over some contractible 4-manifolds $W$ that $Z_{m,n}\#(-Z_{m,n})$ bounds, then we would obtain a simply-connected, equivariant definite cobordism $W^{\prime}$ from $(Z_{m,n},\tau)$ to itself.
+As in the proof of Theorem 1.2, $r_{0}(Z_{m,n},\tau)<\infty$.
+By Theorem 3.7, $r_{0}(Z_{m,n},\tau)<r_{0}(Z_{m,n},\tau)$, which is a contradiction.
 Proof of Theorem 1.6.
@@
 The same holds for the involution $\sigma$. ∎
-Figure 19. The simply-connected, equivariant negative-definite cobordism from $(S^{3}_{1/{m+1}}(K_{n}),\tau)$ to $(S^{3}_{1/m}(K_{n}),\tau)$ used in the proof of Theorem 1.5.
+Figure 19.
+The simply-connected, equivariant negative-definite cobordism from $(S^{3}_{1/{m+1}}(K_{n}),\tau)$ to $(S^{3}_{1/m}(K_{n}),\tau)$ used in the proof of Theorem 1.6, obtained by attaching (1)(-1)-framed 2-handles along the green unknot.
 Remark 3.13.